OpenStudy (anonymous):

Solve x – 4 = –7. A. x = 3 B. x = –3 C. x = 11 D. x = –11

OpenStudy (anonymous):

subtract 4 brom both sides

OpenStudy (anonymous):

\[x-4=-7\]\[x-4+4=-7+4\]\[x=-3\] Check: \[x-4=-7\]\[(-3)-4=-7\]\[-7=-7\] CHECK